Taylor put $200 into a CD that pays 6% interest, compounded quarterly. According to the rule of 72, approximately how long will it take for his money to double?
Answer by Alan3354(69443)

Taylor put $200 into a CD that pays 6% interest, compounded quarterly. According to the rule of 72, approximately how long will it take for his money to double?
---------------
400 = 200*(1 + 0.06/4)^t t = quarters
(1 + 0.015)^t = 2
t*ln(1.015) = ln(2)
t = ln(2)/ln(1.015)
t =~ 46.555 quarters
========================
I don't know the rule of 72.
If you know it, you can do it that way.